### v1

@@ 3. Model
We consider a network of $n\geq 3f+2p+1$ participants called replicas. Our protocol is safe and live with up to $f$ replicas being Byzantine. Additionally, we guarantee fast termination when up to $p$ replicas are not cooperating. In other words, the parameter $p\in[0,f]$ denotes the maximum number of replicas that are not needed for the fast path to be successful.
In particular, by setting $p=0$, we reach the upper bound on the number of Byzantine replicas permissible (Dwork et al., 1988), i.e. $n\geq 3f+1$. In this case, the fast path will be used only when all replicas, including Byzantine ones, are in pre-agreement on a block and behaving honestly. On the other side of the spectrum, the fast path can be made robust against Byzantine behavior, by setting $p=f$.
We consider the same communication model used in ICC, in which replicas communicate over a partially synchronous network (Dwork et al., 1988). Partial synchrony lies between the definitions of a synchronous network and an asynchronous network. In a synchronous network, there is a known fixed upper bound $\Delta_{bnd}$ (or simply $\Delta$) on the time required for a message to be sent from one party to another. In an asynchronous network no fixed upper bounds $\Delta_{bnd}$ exists. We use $\delta$ to denote the true message delivery time of replicas, i.e., the unknown time it takes for one communication step. In partial synchrony, the network alternates intervals of synchrony, in which the bound $\delta<\Delta_{bnd}$ holds, to ones of asynchrony.
Replicas communicate via all-to-all authenticated links. We assume the existence of a public-key infrastructure (PKI), secure digital signatures, and collision-resistant hash functions. Additionally, a random leader election oracle is assumed to exist.
@@ 4. Internet Computer Consensus
The Internet Computer Consensus protocol (ICC) is used to make replicas in the same network agree on a chain of blocks. As the protocol progresses, a tree of blocks is constructed, starting from a genesis block that is at the root of the tree. While each replica may have a different, partial view of the tree, no replicas have a conflicting view of the tree. As the replicas make progress, by finalizing at most one block per tree height, a path of finalized blocks is established. This path constitutes the blockchain. Honest replicas have a consistent view of this path.
Creating the Block-Tree through Notarization. The ICC protocol proceeds in rounds. In the k-th round, one or more blocks are added to the tree at height $k$. In the following, we will use round and height interchangeably. To be added to the tree, a block must be notarized. In ICC, a block $b$ becomes notarized for a replica, once at least $n-f$ notarization shares for block $b$ are received. A notarization share is a signature share of a BLS multi-signature scheme (Boneh et al., 2018) in which a signature can be reconstructed by combining any $n-f$ of the $n$ signature shares, each signed by a different replica. A notarization share sent by a replica $u$ for a block $b$ implies that replica $u$ validated block $b$. As soon as block $b$ is validated by at least $n-f$ replicas, and thus becomes notarized, it can be added to the block-tree.
Block Proposal. In principle, in every round $k$, each replica can propose a block and all notarized blocks make it into the block-tree. Therefore, at any height $k$ there can be up to $n$ blocks. However, this makes it difficult for the replicas to agree on which block should be part of the blockchain, as at most one block per height should be included. To reduce the number of blocks proposed in each round, a random beacon is used to generate a random permutation of the $n$ replicas. The permutation defines a different rank $r\in[0,n-1]$ for each replica for that particular round. The replica with the lowest rank, i.e., the rank-0 replica, is the leader of that round. At the beginning of a round, each replica computes its rank and starts a timer. The leader of that round proposes the block immediately, while the other replicas wait for a time directly proportional to their rank before proposing a block. Let $r$ be the rank of a replica in a given round, the replica delays the proposal of its block by a proposal delay $\Delta_{prop}(r)=2\Delta_{bnd}r$. In round $k$, when deciding which block in the block-tree to extend, the replica considers only notarized blocks at height $k-1$.
Notarization. In round $k$, before sending a notarization share for a block $b$ proposed by a replica $u$ with rank $r^{(u)}$, the replica $v$ receiving the block $b$ waits a notarization delay of $\Delta_{notary}^{(u)}=2\Delta_{bnd}r^{(u)}$ after starting round $k$. Therefore, once replica $v$ receives a block $b$ from replica $u$, it checks if it is time to send a notarization share for a block of rank $r^{(u)}$. If so, it broadcasts the notarization share to all other replicas, otherwise, it waits until $\Delta_{notary}^{(u)}$ has passed. This way, blocks of lower rank should become notarized and added to the block-tree before others. Once a block $b$ in round $k$ becomes notarized for replica $u$, i.e., replica $u$ receives at least $n-f$ notarization shares for $b$ and combines them into a notarization, replica $u$ broadcasts the notarization to other replicas in the subnet, stops sending notarization shares for blocks of round $k$, and starts round $k+1$. Thus, when the leader is honest, the network is in a phase of synchrony and the block proposal and notarization delays are set accordingly, only the block proposed by the rank-0 replica will be added to the tree at height $k$. If the leader is not honest or the network is asynchronous, some other replicas of higher rank may also propose blocks, and also have their blocks added to the tree.
Finalization. As there might be multiple notarized blocks at the same height, we must guarantee that all replicas agree on the same finalized block and that all finalized blocks are part of the same blockchain. Each replica $u$ could assume that after a certain time period, it will have seen all notarized blocks that have been proposed, and if there is just one notarized block at a height $k$, it considers the chain to be agreed upon up to this height. However, this approach does not provide a deterministic finalization guarantee as perhaps the network is in a moment of asynchrony and there are other notarized blocks at height $k$ that the replica $u$ has not received yet. If the replica does not wait long enough, it may violate safety. Thus, on one side blocks need to be agreed upon quickly to reduce the block finalization delay as much as possible, but on the other side, replicas might have to wait a long time to not break consistency. To avoid this trade-off, ICC uses a mechanism similar to notarization to guarantee deterministic finalization. If the replica did not previously broadcast notarization shares for any other block in the same round $k$, the replica will also broadcast another signature share for block $b$, called finalization share. If a replica $u$ receives at least $n-f$ finalization shares for the same block $b$, the replica $u$ can combine them into a finalization. At this point, the replica $u$ is guaranteed that all replicas will eventually include block $b$ in the blockchain. Once a finalization is created for a given block $b$, block $b$ is said to be explicitly finalized and can be output. The finalization is then sent to all the other replicas in the subnet.
The rule for broadcasting finalization shares implies that in some rounds no block might become finalized by the aggregation of at least $n-f$ finalization shares. However, as soon as a block in a later round is explicitly finalized, all its ancestors in the tree, until the last explicitly finalized block, automatically become implicitly finalized. An illustration of the steps necessary before reaching finalization is shown in fig. 3.
@@
Figure 3. The diagram shows the steps necessary to reach the finalization of a block. Black blocks are finalized, dark gray blocks are notarized, while light gray blocks are not. In this example, $n=4$ and $f=1$. Initially, $n-f=3$ replicas send notarization shares (NS) for the rank-0 block. Replicas will not send a NS for higher rank proposals, in this round. As soon as the NS are received, the replica combines them in a notarization (N). Then, the replicas that only sent an NS for the rank-0 block will also send a finalization share (FS) for it. Once the three FS are received, they will be combined into a finalization (F) and both the r-0 block and its parent(s) are finalized.
Lemma 4.1 ().
Let $n\geq 3f+1$, and $b$ be a finalized block at height $k$. ICC guarantees that if block $b^{\prime}$ is notarized at height $k$, then $b=b^{\prime}$.
Proof.
Let a block $b$ at height $k$ be finalized for replica $u$. Replica $u$ must have received at least $n-f$ finalization shares for block $b$, $n-2f$ of which are from honest replicas. Each honest replica sends a finalization share for a notarized block $b$ at height $k$ only if it did not send any other notarization share for other blocks at height $k$. Therefore, other blocks at height $k$ receive at most $2f$ notarization shares. The bound $n>3f$ implies that $2f<n-f$, thus other blocks at height $k$ do not receive enough notarization shares to become notarized. ∎
@@
In other words, if a block $b$ at height $k$ becomes finalized, no other block at height $k$ can be notarized. Since only notarized blocks are extended, lemma 4.1 implies that all blocks at height greater than $k$ will have $b$ as an ancestor and thus, all finalized blocks are part of the same blockchain. This sketches the safety proof of the ICC algorithm.
@@
[Finalization Latency] In rounds in which the network is synchronous, finalization of a block can be reached in three times the message delivery time $\delta<\Delta_{bnd}$. In the first round the block proposal, in the second the notarization shares, and in the third the finalization shares are broadcast (see fig. 1).
@@
[Timeouts] In the simplest implementation of the ICC protocol, we can assume that the communication delay bound $\Delta_{bnd}$ is an explicit parameter. In practice, instead, the protocol is modified to adaptively adjust to an unknown communication delay bound.
@@ 5. Problem Statement
The role of the consensus layer of the Internet Computer is to order blocks containing an arbitrary payload so that all replicas output the payload in the same order. The properties we want the Fast Internet Computer Consensus algorithm to satisfy, are first and foremost the same that Internet Computer Consensus provides (Camenisch et al., 2022):
  • Deadlock Freeness: In each round, the block-tree height increases by 1.
@@
Additionally, Fast Internet Computer Consensus satisfies a stronger property:
@@
Property 5 is also a liveness property. It provides chain-growth even in periods of asynchrony, while Property 5 guarantees that the entire progress is finalized in periods of synchrony. This distinction displays the two modes of operation (Camenisch et al., 2022).
@@ 6. Intuition
The goal of the Fast Internet Computer Consensus fast path is to finalize a block as soon as possible while guaranteeing consensus among replicas. In the ICC protocol, a block is finalized at the fastest three communication steps after being proposed. The fast path added in FICC reduces the finalization latency down to two communication steps, as shown in fig. 1.
Definition 6.1 ().
@@
In order to distinguish the two possible finalization scenarios, we call Fast Path finalized (or FP-finalized) the blocks that are explicitly finalized via the FICC fast path, while we call Internet Computer finalized (or IC-finalized) the blocks that are explicitly finalized by receiving at least $n-f$ finalization shares, as in the ICC protocol.
@@
The idea behind the fast path is that if “enough” replicas send their first notarization shares for the same block $b$, i.e., replicas are in pre-agreement, then $b$ can be immediately finalized without waiting for the finalization shares to be sent and received. In other words, replicas fall back to using the finalization through finalization shares only if there are “too many” replicas that do not agree on the first block added to the block-tree at a given height.
Definition 6.2 (Fast Share).
@@
To determine if enough replicas are in pre-agreement, and try to finalize a block via the fast path, replicas broadcast a fast share for the first block they notarize in each round. A block that receives at least $n-p$ fast shares becomes explicitly finalized via the fast path, and fast shares can be combined into a fast finalization.
@@
Safety considerations. As we have sketched in lemma 4.1, the rules of ICC ensure that if a block $b$ at height $k$ becomes IC-finalized, no other block at the same height will be notarized. This guarantees that the tree will not contain any block at height $k$ besides block $b$. Therefore, only block $b$ will be extended. In FICC however, if a block $b$ gets FP-finalized, we cannot guarantee that there will not be other notarized blocks at the same height. If implemented without care, this could induce the finalization of conflicting blocks.
@@
Hence, we introduce a new concept, similar to notarization. Intuitively, we want to guarantee that whenever a block $b$ at height $k$ gets FP-finalized, then no other block at height $k$ can be extended. To this end, each replica marks blocks as fastable when they are safe to be extended.
@@
As we will show, the conditions of blocks being fastable almost never restrict the original ICC algorithm, and only a few additions are sufficient to guarantee the safety of FICC.
7. Fast Internet Computer Consensus
@@
The FICC algorithm is an extension to the Internet Computer Consensus algorithm that enables fast path finalization in periods of synchrony. The protocol provides additional conditions that, if satisfied, guarantee explicit finalization to be achieved as soon as enough fast shares are received. The fast path runs alongside the Internet Computer Consensus protocol and is integrated into existing messages. In case the fast path cannot be used, there is no “switching cost” to revert to the slow path. When there is no pre-agreement, blocks will still be notarized and eventually finalized (either explicitly or implicitly) just as in the ICC protocol.
@@
The following definitions are written from the point of view of a replica $u$:
Definition 7.1 ().
Let B(k) be the set of blocks that have been received in round $k$. Let FS(b) be the set of replicas that have sent a fast share for $b\in B(k)$.
Definition 7.2 ().
At any given time, m(k) is the round-k block for which the most fast shares have been observed:
$m(k)=\operatorname*{arg\,max}_{b\in B(k)}{\left|FS(b)\right|}$
Definition 7.3 ().
At any given time, sum(k) is the number of distinct replicas that have sent a fast share for at least one round-k block:
$sum(k)={\left|\bigcup_{b\in B(k)}FS(b)\right|}$
@@
Using these definitions, we can pin down exactly what blocks are safe to extend for a replica $u$, in the presence of the fast path. We call such a block fastable.
Definition 7.4 (Fastable Blocks).
@@
Finalized blocks are fastable by definition. Additionally, at any point during a round $k$:
@@
    if ${\left|FS(b)\right|}>f+p$, then $b$ is fastable.
@@
    if $sum(k)-{\left|FS(m(k))\right|}>f+p$, all current and future blocks of round $k$ are fastable.
An example of the above definitions is presented in fig. 4.
Figure 4. Fast shares (FaS) received for each block are shown in the block-tree. Black blocks are finalized, dark gray blocks are fastable, while light gray blocks are not. Supposing $n=4$, $f=1$, and $p=0$, for round $k$, as $sum(k)=3$, only the first block is fastable. Instead, for round $k+1$, $sum(k)=4$ and as $4-2>1$, all blocks in this round are fastable.
The Fast Internet Computer Consensus algorithm FICC is defined by the following changes to the ICC algorithm presented in section 4:

### v2

@@ 3. Model
We consider a network of $n\geq\max{(3f+2p-1,3f+1)}$ participants called replicas. Our protocol is safe and live with up to $f$ replicas being Byzantine. Additionally, we guarantee fast termination when up to $p$ replicas are not cooperating. In other words, the parameter $p\in[0,f]$ denotes the maximum number of replicas that are not needed for the fast path to be successful. Note that there is no reason to choose $p=0$, as our protocol will be strictly faster with $p=1$, and require the same number of replicas. We can thus also write $n\geq 3f+2p^{*}-1$, where $p^{*}\in[1,f]$.
By setting $p=1$, we reach the upper bound on the number of Byzantine replicas permissible (Dwork et al., 1988), i.e. $n\geq 3f+1$. In this case, the fast path will be used when the leader and all but one replica behave honestly. On the other side of the spectrum, given an honest leader, the fast path can be made robust against Byzantine behavior, by setting $p=f$.
We consider the same communication model used in ICC, in which replicas communicate over a partially synchronous network (Dwork et al., 1988). In a synchronous network, there is a known fixed upper bound $\Delta$ on the time required for a message to be sent from one party to another. In an asynchronous network, no fixed upper bound $\Delta$ exists. We use $\delta$ to denote the true message delivery time of replicas, i.e., the unknown time it takes for one communication step across all replicas. In partial synchrony, the network alternates intervals of synchrony, in which the bound $\delta<\Delta$ holds, to ones of asynchrony.
Replicas communicate via all-to-all authenticated links. We assume the existence of a public-key infrastructure (PKI), secure digital signatures, and collision-resistant hash functions. Additionally, we assume replicas have access to shared randomness, e.g., through a safe and live random beacon protocol (Raikwar and Gligoroski, 2022).
@@ 4. Internet Computer Consensus
The slow path of the Banyan algorithm corresponds almost precisely to the Internet Computer Consensus protocol (ICC), which we thus introduce first.
@@
Figure 3. The diagram shows the steps necessary to reach the finalization of a block. Black blocks are finalized, dark gray blocks are notarized, while light gray blocks are not. In this example, $n=4$ and $f=1$. Initially, $n-f=3$ replicas send notarization votes (NV) for the rank-0 block of round $k$. Replicas will not send a NV for higher rank proposals, in this round. As soon as the NV are received, the replica combines them in a notarization (N). Then, the replicas that only sent an NV for the rank-0 block will also send a finalization vote (FV) for it. Once the three FV are received, they will be combined into a finalization (F) and both the r-0 block and its ancestor(s) are finalized.
Overview. The goal of the ICC protocol is to provide a total ordering (i.e., chain) of blocks among replicas in a network. As the protocol advances, a tree of blocks is constructed, starting from a genesis block that is at the root of the tree. Blocks are added to the block-tree if they are safe to be extended. Each replica may have a different, partial view of the block-tree. Replicas make progress, by finalizing at most one block per tree height. By traversing the tree edges, a path of finalized blocks is established. This path constitutes the blockchain. Honest replicas are guaranteed to have a common prefix of this path.
Creating a Block-Tree through Notarization. The ICC protocol proceeds in rounds. In the k-th round, one or more blocks are added to a block-tree at height $k$. To be added to the tree, a block must be notarized. In ICC, a block $b$ becomes notarized for a replica, once at least $n-f$ notarization votes for block $b$ are received. A notarization vote is a BLS signature sent by a replica $u$ for a block $b$, implying that replica $u$ validated block $b$. Notarization votes can be aggregated to a single multi-signature, that can be efficiently verified (Boneh et al., 2018). As soon as block $b$ is validated by at least $n-f$ replicas, it becomes notarized, and can be added to the block-tree.
Block Proposal. In principle, in every round $k$, each replica can propose a block and all notarized blocks make it into the block-tree. Therefore, at any height $k$ there can be up to $n$ blocks. However, this makes it difficult for the replicas to agree on which block should be part of the blockchain, as at most one block per height should be included. To reduce the number of blocks proposed in each round, a random beacon is used to generate a random permutation of the $n$ replicas. The permutation defines a different rank $r\in[0,n-1]$ for each replica for that particular round. The replica with the lowest rank, i.e., the rank-0 replica, is the leader of that round. At the beginning of a round, each replica computes its rank and starts a timer. The leader of that round proposes the block immediately, while the other replicas wait for a time directly proportional to their rank before proposing a block. Let $r$ be the rank of a replica in a given round, the replica delays the proposal of its block by a proposal delay $\Delta_{prop}(r)=2\Delta\times r$. In round $k$, when deciding which block in the block-tree to extend, the replica considers only notarized blocks at height $k-1$.
Notarization. In round $k$, before sending a notarization vote for a block $b$ proposed by a replica $u$ with rank $r^{(u)}$, the replica $v$ receiving the block $b$ waits a notarization delay of $\Delta_{notary}^{(u)}=2\Delta\times r^{(u)}$ after starting round $k$. Therefore, once replica $v$ receives a block $b$ from replica $u$, it checks if it is time to send a notarization vote for a block of rank $r^{(u)}$. If so, it broadcasts the notarization vote to all other replicas, otherwise, it waits until $\Delta_{notary}^{(u)}$ has passed. This way, blocks of lower rank should become notarized and added to the block-tree before others. Once a block $b$ in round $k$ becomes notarized for replica $u$, replica $u$ broadcasts the notarization to other replicas, stops sending notarization votes for blocks of round $k$, and starts round $k+1$. Thus, when the leader is honest, the network is in a phase of synchrony and the block proposal and notarization delays are set accordingly, only the block proposed by the rank-0 replica will be added to the tree at height $k$. If the leader is not honest or the network is asynchronous, some other replicas of higher rank may also propose blocks, and also have their blocks added to the tree.
Finalization. As there might be multiple notarized blocks at the same height, we must guarantee that all replicas agree on the same finalized block and that all finalized blocks are part of the same blockchain. ICC uses a mechanism similar to notarization to guarantee that no consistency violation across rounds is possible. If the replica did not broadcast notarization votes for any other block than $b$ in the same round $k$, the replica will also broadcast another BLS-signature for block $b$, called finalization vote. If a replica $u$ receives at least $n-f$ finalization votes for the same block $b$, the replica $u$ can aggregate them into a finalization. At this point, the replica $u$ is guaranteed that all replicas will eventually include block $b$ in the blockchain. Once a finalization is created for a given block $b$, block $b$ is said to be explicitly finalized and can be output. The finalization is then sent to all the other replicas.
The rule for broadcasting finalization votes implies that in some rounds, no block might become finalized by the aggregation of at least $n-f$ finalization votes. However, as soon as a block in a later round is explicitly finalized, all its ancestors in the tree, until the last explicitly finalized block, automatically become implicitly finalized. An illustration of the steps necessary before reaching finalization is shown in Figure 3.
@@
We point out that if a block $b$ at height $k$ becomes finalized, no other block at height $k$ can be notarized. Since only notarized blocks are extended (see beginning of Section 4), this implies that all blocks at height greater than $k$ will have $b$ as an ancestor and thus, all finalized blocks are part of the same blockchain. This sketches the safety proof of the ICC algorithm.
@@
[Finalization Latency] In rounds in which the network is synchronous, finalization of a block can be reached in three times the message delivery time $\delta<\Delta$. In the first round the block proposal, in the second the notarization votes, and in the third the finalization votes are broadcast (see Figure 1).
@@
[Timeouts] In the simplest implementation of the ICC protocol, we can assume that the communication delay bound $\Delta$ is an explicit parameter. In practice, instead, the protocol is modified to adaptively adjust to an unknown communication delay bound.
@@ 5. Problem Statement
The purpose of SMR protocols is to totally order blocks containing an arbitrary payload so that all replicas output the payload in the same order.
The properties we want the Banyan algorithm to satisfy, are first and foremost the same that Internet Computer Consensus provides (Camenisch et al., 2022):
  • Deadlock Freeness: Each round eventually terminates and increases the block-tree height by 1.
@@
Additionally, Banyan satisfies a stronger property:
@@
Deadlock freeness is also a liveness property. It provides chain growth even in periods of asynchrony, while Liveness guarantees that the entire progress is finalized in periods of synchrony. This distinction underlines the two modes of operation (Camenisch et al., 2022).
@@ 6. Intuition
The goal of the Banyan fast path is to finalize a block as soon as possible while guaranteeing agreement among replicas. In the ICC protocol a block is finalized at the fastest three communication steps after being proposed. The fast path added in Banyan reduces the finalization latency down to two communication steps, as shown in Figure 1.
Definition 6.1.
@@
In order to distinguish the two possible finalization scenarios, we call Fast Path finalized (or FP-finalized) the blocks that are explicitly finalized via the Banyan fast path, while we call Slow Path finalized (or SP-finalized) the blocks that are explicitly finalized by receiving at least $n-f$ finalization votes, as in the ICC protocol.
@@
The idea behind the fast path is that if “enough” replicas send their first notarization votes for the same block $b$, i.e., replicas are in pre-agreement, then $b$ can be immediately finalized without waiting for the finalization votes to be sent and received. In other words, replicas finalize blocks through finalization votes only if there are too many replicas that do not agree on the first block added to the block-tree at a given height, or if the slow path is faster.
Definition 6.2 (Fast Vote).
@@
To determine if enough replicas are in pre-agreement, and try to finalize a block via the fast path, replicas broadcast a fast vote for each round’s first block they send a notarization vote for. A block that receives $n-p$ fast votes becomes explicitly finalized via the fast path, and fast votes can be combined into a fast finalization.
@@
Safety considerations. As we have sketched in Section 4, the rules of ICC ensure that if a block $b$ at height $k$ becomes SP-finalized, no other block at the same height will be notarized. This guarantees that the tree will not contain any block at height $k$ besides block $b$. Therefore, only block $b$ will be extended. In Banyan however, if a block $b$ gets FP-finalized, we cannot guarantee that there will not be another notarized blocks at the same height.
@@
Hence, we introduce a new concept, similar to notarization. Intuitively, we want to guarantee that whenever a block $b$ at height $k$ gets FP-finalized, then no other block at height $k$ can be extended. To this end, each replica marks blocks as unlocked when they are safe to be extended.
@@
As we will show, the conditions of a block being unlocked almost never restrict the original ICC algorithm (see Section 8.1) and only a few additions are sufficient to guarantee the safety of Banyan.
7. The Banyan Algorithm
@@
The Banyan algorithm extends the Internet Computer Consensus algorithm and enables fast path finalization in periods of synchrony. The protocol allows explicit finalization to be achieved as soon as $n-p$ fast votes are received. The fast path runs alongside the Internet Computer Consensus protocol and is integrated into existing messages. In case the fast path cannot be used, there is no “switching cost” to revert to the slow path. When there is no pre-agreement, blocks will still be notarized and eventually finalized (either explicitly or implicitly) just as in the ICC protocol.
@@
The following definitions are written from the point of view of a replica $u$, at round $k$:
Definition 7.1.
Let $\textsc{blocks}(k)$ be the set of blocks that have been received in round $k$. For a block $b\in\textsc{blocks}(k)$, we define $\textsc{supp}(b)$ to be the set of replicas from which $u$ has received a fast vote. Further, for any set $\mathcal{B}\subseteq\textsc{blocks}(k)$, we define $\mathbf{\textsc{supp}(\mathcal{B})}$ to be the set of distinct replicas from which $u$ has received a fast vote for a block in $\mathcal{B}$:
$\textsc{supp}(\mathcal{B})=\bigcup_{b\in\mathcal{B}}\textsc{supp}(b)$
Definition 7.2.
At any given time, $\textsc{max}(k)$ is a rank-0 block, such that no other rank-0 block has a larger support:
$\textsc{max}(k)=\left(\operatorname*{arg\,max}_{b\in\textsc{blocks}(k),b.rank=0}{\left|\textsc{supp}(b)\right|}\right).pop()$
{remark}
There can be multiple rank-0 blocks when the leader is Byzantine.
Definition 7.3.
At any given time, $\textsc{nonLeaderBlocks}(k)$ is the set of blocks which $u$ has received with a rank larger than 0:
$\textsc{nonLeaderBlocks}(k)=\{b\in\textsc{blocks}(k)\mid b.rank\neq 0\}$
Definition 7.4.
At any given time, $\textsc{nonMaxBlocks}(k)$ is the set of blocks which $u$ has received, excluding a rank-0 block that has the most support:
$\textsc{nonMaxBlocks}(k)=\{b\in\textsc{blocks}(k)\mid b\neq\textsc{max}(k)\}$
@@
Using these definitions, we can pin down exactly what blocks are safe to extend for a replica $u$, in the presence of the fast path. We call such a block unlocked.
Definition 7.5 (Unlocked Blocks).
@@
Finalized blocks are unlocked by definition. Additionally, at any point during a round $k$, for a valid block $b\in\textsc{blocks}(k)$:
@@
    if ${\left|\textsc{supp}(b)\right|}+{\left|\textsc{supp}(\textsc{nonLeaderBlocks}(k))\right|}>f+p$, then $b$ is unlocked.
@@
    if ${\left|\textsc{supp}(\textsc{nonMaxBlocks}(k))\right|}>f+p$, all current and future blocks of round $k$ are unlocked.
An example of the above definitions is presented in Figure 4.
Figure 4. Fast votes (FaV) received for each block are shown in the block-tree. Black blocks are finalized, dark gray blocks are unlocked, while light gray blocks are not. Supposing $n=4$, $f=1$, and $p=1$, for round $k$, Condition 1 is met, and the r-0 block is unlocked. Instead, for round $(k+1)$, Condition 2 is met, and all blocks in this round are unlocked.
The Banyan algorithm is defined by the following changes to the slow path algorithm (ICC) presented in Section 4. Line numbers refer to Algorithms 1 and 2 that contain the full Banyan pseudocode.